Find the surface area of a cylinder. ROUND TO THE NEAREST TENTH.
1.) r= 6.5 m , h= 15 m

2.) d= 8.7 in , h= 7.3 in

Answer by checkley75(3666)

1)TSA=(2pi*r*h)+(2pir^2)
TSA=(2*3.14*6.5*15)+(2*3.14*6.5^2)
TSA=2112.3+265.33
TSA=2377.63 m^2
--------------------------------
2)TSA=(pi*d*h)+(2pir^2)
TSA=(3.14*8.7*7.3)+(2*3.14*7.3^2)
TSA=199.42+334.66
TSA=534.08 in^2